Entered Answer Preview Result 0.3708 0.3708 correct 0.6476 0.6476 Incorrect At least one of the answers above is NOT correct. 1 point) The probability of the closing of the ith relay in the circuits shown is given by Pi. Let py-0.8, P2 0.5, Ps-0.2,P1-0.1,Ps-0.9. If all relays function Independently. what is the probablity that a current flows between A and B for the respective circults? (a) P0.3708 (b) P 0.6476
Answer #1

a)

P(current flow) =(1-P(1-2 not works)*(3-4 not works))*5 works=(1-(1-0.8*0.5)*(1-0.2*0.1))*0.9=0.4120*0.9

=0.3708

b)

for current to flow from A to B:

p1p4+p1p3p5+p2p5+p2p3p4-p1p3p4p5-p1p2p4p5-p1p2p3p4-p1p2p3p5-p2p3p4p5+2p1p2p3p4p5=0.5590